OpenStudy (aravindg):

Applying laplace equation:

OpenStudy (aravindg):

h=f(y-z),f(z-x),f(x-y) Prove that sum of partial derivatives of h with respect to x,y,z=0

OpenStudy (austinl):

So we are wanting to prove this... \(\displaystyle \frac{\partial h}{\partial x}=\frac{\partial}{\partial x}(f(y-z))=0\) And we have to prove this using The Laplace Transform?

OpenStudy (aravindg):

Laplace equation I guess. I am new to this section of questions.

OpenStudy (anonymous):

not Laplace Transform

OpenStudy (austinl):

Do you have any clue as to how you would set this up? I am not having anything jump out at me.

OpenStudy (aravindg):

We need to show